OpenStudy (halfdeafdarling):

Okay! I have a second part to a question I am looking for. Anyone able to take about 10 to 20 minutes explaining something (may take less time not sure) I really need help on figuring this out so far got the best help from @peachpi thank you so much, hope someone can help just as much!

Directrix (directrix):

What is the question?

OpenStudy (halfdeafdarling):

Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x².

OpenStudy (halfdeafdarling):

This is the second piece I need help on i have the first part done already, but with this im a little confused on what they want

OpenStudy (halfdeafdarling):

@FibonacciChick666

OpenStudy (halfdeafdarling):

took me way too long to figure out how to tag people xD

OpenStudy (fibonaccichick666):

ok, so you could simplify x^3 x^4 right?

OpenStudy (halfdeafdarling):

well wouldnt that be like 1/x^3 being 1/x*x*x?

OpenStudy (halfdeafdarling):

or am i wrong with that?

OpenStudy (fibonaccichick666):

well, that is one way to rewrite x^(-3)

OpenStudy (fibonaccichick666):

so, all we are doing is ending up with a 2 instead of a 3

OpenStudy (halfdeafdarling):

okay tbh though I am still a little confused >.<

OpenStudy (halfdeafdarling):

well actually I have another part I might just have a question on seeing if it is correct. Substitute your birthday into the equation Square root of the quantity x minus y + m = d, where y is the last two digits of your birth year, m is the month, and d is the day. If you were born on 7/10/1856 like Nikola Tesla, your equation would be Square root of the quantity x minus 56 + 7 = 10. Solve for x and identify if it is an extraneous solution. I have as my equation \[\sqrt{x-0} +1 = 4\]

OpenStudy (halfdeafdarling):

for this i then begun subtracting 1 getting 3

OpenStudy (halfdeafdarling):

once I subtracted that I had squared the sides giving me then x-0=9 thus x equalling 9. Once I had that i went on with \[\sqrt{9}\] which equals 3. thus being 3-0+1=4 being that it is correct?

Directrix (directrix):

@HalfDeafDarling Did you abandon this project: >>>equivalent forms of x²

OpenStudy (halfdeafdarling):

no i will still be going over that and wish for help on it, but the one I just did might seem like an easier one for ask for now if i even have it done correctly since that is the one I have done fully already

OpenStudy (fibonaccichick666):

I'm sorry, I can't help. I just got called to do something. Direct can help you through

OpenStudy (halfdeafdarling):

I ACTUALLY GOT IT THANKS YALL

Directrix (directrix):

Happy to hear that.